two trains 75 km apart from each other on parallel tracks each moving at 75 mph. a bird flies back and forth between the trains at 20 kmph until the train pass each other . how far does the bird fly

Grade:11

2 Answers

Hadiya
6 years ago
Let the trains meet in t time. Total distance equal to 75 km.Total distance equal to distance travelled by first train + distance travelled by second train.Distance equal to speed ×time.75t÷1000+75t÷1000=75km.t=5000hDistance travelled by the bird =speed of the bird × time time in which the trains meet.=20×500=1000km.
Luan Phan
2 years ago
The train 75 km apart approach each other en parallel tracks, each moving at
kmh. A bird flies back and forth between the trains at 20km/h until the trains pass
each other. How for does the bird fly